OpenStudy (anonymous):

Use square root property to solve the equation: x^2=20 i got sqrt2? is this right?

OpenStudy (anonymous):

\[x^2=20\] \[x=\pm\sqrt{20}\] \[x=\pm\sqrt{4\cdot5}\] \[x=\pm2\sqrt{5}\]

OpenStudy (anonymous):

awesome! XD i got that the first time then i thought it looked dumb and left out the five, thanks!!! :)

OpenStudy (anonymous):

if it's not perfect square like 2*2=4 3*3=9 it will look dumb :D

OpenStudy (anonymous):

^^ haha yeah